OpenStudy (anonymous):

Use 5 step problem solving method. Angle G and Angle H are supplementary angles. If Angle H is ten less than four times Angle G, what are the measures of the angles?

OpenStudy (anonymous):

Let angleG = x degree then angleH = 4(angleH) - 10 = 4*x - 10 = 4x -10 Because the angles are supplementary, it means their sum is 180 deg (x) + (4x - 10) = 180 5x = 180 + 10 x = 190/5 x = 38 Hence angleG = 38 deg and angleH = 4(38) - 10 = 152 - 10 = 142

OpenStudy (anonymous):

THANK YOU!!!!!!!!!

OpenStudy (anonymous):

makes so much sense now!

OpenStudy (anonymous):

u r welcome....☺

OpenStudy (anonymous):

although in the second line would it be: 4(angle g) - 10 ?

OpenStudy (anonymous):

all of it makes perfect sense now except that... ? help? :)